Two sides of a triangle are 3 feet and 6 feet longer than the third side of the triangle, respectively. If the perimeter of the triangle is 36 feet, find the length of each side.
Answer by rfer(16322)

x+(x+3)+(x+6)=36
3x+9=36
3x=27
x=9
x+3=12
x+6=15